OpenStudy (anonymous):

Solve. m^-5 =1/32 m =

OpenStudy (anonymous):

@Vocaloid

OpenStudy (isaiah.feynman):

\[m^{-5} = \frac{ 1 }{ 32 }\]

OpenStudy (anonymous):

exactly

OpenStudy (isaiah.feynman):

Recall that \[m^{-5}=\frac{ 1 }{ m^{5} }\]

OpenStudy (anonymous):

yes. Kind of.

OpenStudy (isaiah.feynman):

So, \[\frac{ 1 }{ m^{5} } = \frac{ 1 }{ 32 }\]

OpenStudy (anonymous):

Do I divide?

OpenStudy (isaiah.feynman):

Do a cross multiplication.

OpenStudy (anonymous):

I am not getting it.

OpenStudy (anonymous):

@Isaiah.Feynman

OpenStudy (isaiah.feynman):

When you cross multiply, you get \[m^{5}=32\]

OpenStudy (isaiah.feynman):

Solve for m from here.

OpenStudy (anonymous):

2?

OpenStudy (isaiah.feynman):

Yes.

OpenStudy (anonymous):

can you help me with a few more?

OpenStudy (isaiah.feynman):

Ummm, one more. Please Open a new thread.

OpenStudy (anonymous):

ok